OpenStudy (anonymous):

Find the product: (1/7x^2 - 2/5)(1/2x-1/2)= (1/7x^2)(1/2x)(1/7x^2)(-1/2)(-2/5)(1/2x)(-2/5)(-1/2) = (1/14x^3)(-1/14x^2)(-1/5x)(1/5) Can someone explain if I have done this correctly?

OpenStudy (zzr0ck3r):

foil first (1/7)x^2 * (1/2)x = (1/14)x^3 outer (1/7)x^2*(-1/2) = -(1/14)x^2 inner (-2/5)*(1/2)x = (-1/5)x last (-2/5)(-1/2) = 1/5 now add them up (1/14)x^3-(1/4)x^2-(1/5)x + 1/5

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

would the answr be x^5?

OpenStudy (zzr0ck3r):

no the answer is the last line I typed

OpenStudy (anonymous):

okay, this is because we cannot go any furtherei that they are not like terms?